# Bridging the Preference Gap between Retrievers and LLMs

### 13 Jan 2024 v1

@@ Abstract
Large Language Models (LLMs) have demonstrated superior results across a wide range of tasks, while retrieval has long been established as an effective means of obtaining task-relevant information for humans. Retrieval-augmented Generation (RAG) are known for their effectiveness in knowledge-intensive tasks by locating relevant information and placing it within the context window of the LLM. However, the relationship between retrievers and LLMs is still under-investigated. Most existing work treats the retriever and the LLM as independent components and leaves a gap between retrieving human-friendly information and assembling a LLM-friendly context. In this work, we examine a novel bridge model, validate the ranking and selection assumptions in retrievers in the context of RAG, and propose a training framework that chains together supervised and reinforcement learning to learn a bridge model. Empirical results demonstrate the effectiveness of our method in both question-answering and personalized generation tasks.
@@ 1 Introduction
Large language models (LLMs) such as GPT-3 Brown et al. (2020) and PaLM2 Anil et al. (2023), have demonstrated impressive performance on a wide range of language tasks. Alongside LLMs, retrieval-augmented generation (RAG), which can retrieve knowledge from an external dataset when needed, have also produced strong results in many knowledge intensive-tasks.
@@
Figure 1: We observe a preference gap when changing the ranking and selection within the LLM’s input context. The experiments are conducted with GTR Ni et al. (2021) as retriever and a fixed Palm2-S as the LLMs. “Top-K” means we select the top-k GTR retrieved passages. Different colors indicate different datasets. The selected passages are used as additional context. We see that selection (Top-1) significantly affects (either positively or negatively) the model’s performance, while ranking (randomized top-5) does not have much impact (the metrics are detailed in Sec. 5.2).
@@
However, despite some extensive RAG studies Khandelwal et al. (2020); Borgeaud et al. (2022); Izacard et al. (2022); Yasunaga et al. (2023), most of them study retrievers and LLMs separately. On one hand, considerable effort has been dedicated to designing user-friendly retrievers, based on the general belief that ranking is paramount, as humans typically read from top to bottom. On the other hand, LLMs exhibit preferences different from humans and yield accurate predictions only when prompts align with these preferences. This discrepancy leads to sub-optimal design in current RAG systems, a phenomenon we term preference gap. This gap manifests in various aspects. For example, the general belief on ranking may not align with LLM’s preferences due to the self-attention mechanism in Transformers, which can focus on any token regardless of its sequential position. Another aspect is selection; humans can disregard irrelevant passages, but it has been shown that LLMs are highly sensitive to irrelevant content Shi et al. (2023a). There are likely more gaps that further diverge from human behavior, such as repetition, which is generally considered detrimental for retrieval systems Xia et al. (2017), but maybe be useful to LLMs for relevance weighting.
@@
We empirically investigate this preference gap, focusing specifically on ranking and selection. As shown in Fig. 1, the performance varied by around 1% when we randomized the top-5 retrieved passages [^1] We note that this is different from Liu et al. (2023), where a “loss in the middle” phenomenon is observed, meaning that the model are better at using the relevant information at the beginning or end of its input context. This is probably due to different number of passages used. In Liu et al. (2023), 20 total documents are employed, whereas we used 5 passages. Regardless, this does not affect the conclusion that there exists a preference gap between the retrievers and LLMs which needs to be bridged.. However, the variation exceeded 5% when selecting the top-1 passage. This indicates the general belief in the importance of ranking in retrieval does not apply to LLMs , indicating that selection could be more crucial. This finding confirms the existence of the preference gap between the retrievers and LLMs, and highlights the importance of bridging this preference gap to enhance RAG. This is a crucial insight that has not been revealed before, which can guide RAG designers towards what should be achieved in the model.
To design a RAG model, existing work has tried to finetune the LLMs to align with the retriever’s preference, or to finetune the retriever to align with the LLM’s preference. However, since they primarily focuses on ranking, their approaches are sub-optimal for bridging the preference gap. Fine-tuning retrievers can only achieve re-ranking, and fine-tuning LLMs is often impractical, as many are only accessible through APIs. To better bridge the gap beyond ranking, we propose a framework called BGM (Bridging the Gap between retrievers and LLMs). We propose to keep the retriever and LLM fixed, and train a bridge model in between. The bridge model aims to transform the retrieved information into a format that LLMs can work with effectively.
Since the preference gap is present in both ranking and selection, and potentially extends to aspects such as repetition, in contrast to the conventional scoring format, we structure the bridge model as a sequence-to-sequence (seq2seq) model. This approach allows it to not only rerank but also dynamically select passages for each query, and potentially employ more advanced strategies like repetition. In a typical RAG, there is no ground truth relevance label for what should be retrieved but only ground truth label for the downstream task. Existing supervised learning (SL) approaches use the supervision provided by the LLM, such as the perplexity of downstream tasks. However, this is ineffective due to sparse supervision as it is nearly impossible to feed all possible retrieved sequences into the LLM to obtain supervision. Even worse, such SL relies on the intermediate relevance label and is not end-to-end training on the downstream tasks. To address these issues, we employed reinforcement learning (RL) on the SL trained bridge model, where the downstream performance metrics are used as the reward and the bridge model are regarded as a policy model. This framework, chaining SL and RL, provides increased supervision from the downstream task. It also offers the model the flexibility to explore more advanced strategies, such as repetition, in forming the optimal passage sequence.
@@
Our experiments reveal that BGM can enhance the performance of various downstream tasks, such as Question Answering (QA) and personalized generation, across a spectrum of datasets, from public QA and amazon reviews to private email conversations. Notably, the modified passages retrieved by BGM surpass the performance of strong retrievers and reranking baseline models. This underscores the significance and promise of the bridging approach in the realm of RAG. In summary, our contributions can be summarized as follows:
@@
  • We empirically establish the existence of the preference gap between retrievers and LLMs, and introduce BGM, which designed to address this preference gap.
@@
  • The evaluation of BGM comprises diverse datasets, including QA and text generation, including both publicly available and personal information sources. The comprehensive evaluation underscores the effectiveness of BGM in bridging the preference gap and improving performance in downstream tasks.
Figure 2: Differing from previous RAG that update either LLMs, retrievers, or both, BGM not only treats the LLM as a black box but also freezes the retrieval. Only a lightweight bridge model (a Seq2seq LM) is tunable in order to adapt the retrieved information to satisfy the LLM’s preference. This makes BGM applicable to large LLMs, which are often served via APIs and also avoids expensive updates to the retriever.
@@ 2 Related Work
Retrieval-augmented Generation (RAG). Augmenting LLMs with relevant information retrieved from various knowledge sources has shown effective in improving performance across numerous NLP tasks, including language modeling Borgeaud et al. (2022); Khandelwal et al. (2020); Shi et al. (2023b), question answering Lewis et al. (2020); Izacard et al. (2022); de Jong et al. (2023); De Jong et al. (2023); Shi et al. (2023b); Guu et al. (2020); Izacard and Grave (2020); Xu et al. (2023), fact versification Lewis et al. (2020) and text generation Lewis et al. (2020). Specifically, RAG utilizes input as a query and comprises two main components: (1) retriever retrieves a set of items from a large corpus. Note that the retrieval units vary across different works, including documents, passages, or even tokens. In this study, we focus on retrieving passages; and (2) LLM incorporate the retrieved passages as additional information, integrating them into the input context for making final predictions.
A fundamental question in this process arises regarding the differing preferences between LLMs and existing retrievers, with LLMs performing optimally only when their preferences are satisfied. Bridging the gap between these differing preferences is crucial. Depending on which components are subject to updates, this challenge can be categorized into three families
Finetuning retrievers and LLMs jointly. This is the most widely used and conventional setting of RAG Izacard et al. (2022); Khandelwal et al. (2020); Wu et al. (2022); Guu et al. (2020); Lewis et al. (2020). However, they are mostly less than 1B so may nor regard as “large” LM. For example, Altas Izacard et al. (2022) finetunes LLM (T5 Raffel et al. (2020a)) and retriever (Contriever Izacard et al. (2021)) jointly by leveraging the LLM to provide supervisory signal to train the retriever. RAG Lewis et al. (2020) uses a tunnable query encoder and DPR Karpukhin et al. (2020) as retriever, BART as LLM, and design an end-to-end training and inference schema to train the query encoder and the LLM.
@@
Finetuning retrievers only. Although above systems have shown improves results, they are not always applicable in practice. Many LLMs such as GPT-3 and Codex are not open-sourced due to commercial considerations and are only available as black-box APIs. Through these APIs, users can send queries and receive responses. Given this constraint, a natural approach is to only update the retrievers to ensure they can retrieve passages that are more compatible with LLMs. REPLUG Shi et al. (2023b) adapts a similar idea as Atlas but fix the LM. RECOMP Xu et al. (2023) trains a compressors to summarize the retrieved document from retriever. However, this family of models is incapable of performing any sample-level selection and can only choose top passages by setting a fixed threshold.
Unlike the existing three families, BGM works on LLM (> 11B) and directly working on bridging the preference gap, without finetuning LLMs or retrievers but a bridge model in between (Fig. 2).
RL for information retrieval. Before the LLM era, RL has been used in information retrieval (IR) Xia et al. (2017); Wei et al. (2017); Zeng et al. (2018); Xu et al. (2020). The core approach was to frame the IR problem as a Markov Decision Process and apply an RL algorithm to solve it. Typically, an IR task would be structured to determine which document to select for a ranking position, using ranking metrics such as DCG as the reward. This approach had advantages, including using the ranking position directly as supervision and optimizing the IR measure directly. However, none of these existing works explored the application of RL in the context of RAG.
In the LLM era, RL has been used in query rewriting for retrieval Wu et al. (2021); Nogueira and Cho (2017); Adolphs et al. (2021), where a black-box retriever is assumed. This is a different problem from RAG. Bacciu et al. (2023) suggest using RL to fine-tune retriever in the RAG context in their opinion paper, but it does not conduct any experiments. More crucially, their work fails to recognize the importance of bridging the gap between retrievers and LLMs, nor does it specify what the bridge model should be.
@@ 3 Problem Formulation
Retriever. Given an input $x$, the retriever aims to retrieve a ranked list of passages from a corpus $D=\{d_{i}\}_{i=1}^{m}$ that are relevant to $x$. In this work, we assume a frozen dense retriever. Typically, a dual encoder architecture is applied, where an encoder is used to encode both the input context $x$ and the passage $d$. Specifically, the encoder maps each passage to an embedding $\bm{E}(d)$. The similarity between input and passage embedding is computed by their cosine similarity,
@@
The top-k passages that have the highest similarity scores when compared with the input $x$ are retrieved in this step.
@@
Bridge Model for RAG. The retrieved top-K passages provide richer information about the original input $x$ and can potentially help the LLM to make a better prediction on downstream tasks. However, since there is preference gap, we propose a bridge model $\bm{B}$ to adapt the retrieved passages to LLM-friendly passages. As mentioned in the Sec. 1, the bridge model is a seq2seq model. It takes as input all the retrieved passages $(d^{\text{retr.}}_{j})_{j=1}^{k}$ and output a subset of adapted passages $(d^{\text{bdr.}}_{j})_{j=1}^{n}$. This format is advantageous because the seq2seq model automatically considers ranking by generating the next token based on the preceding one and selection by placing the end-of-sentence token in the appropriate position. Note that $n$ may or may not equal to $k$ due to selection.
$(d^{\text{bdr.}}_{j})_{j=1}^{n}=\bm{B}((d^{\text{retr.}}_{j})_{j=1}^{k}).$ (3)
It is important to note that in practice, the bridge model’s input includes the query, passage IDs, and passage content, while the output consists of passage IDs. We then convert the obtained passage IDs to their corresponding documents when we use them in RL (Sec. 4.2). Additionally, the passages in passage sequence (e.g., $(d^{\text{retr.}}_{j})_{j=1}^{n}$) are concatenated into a long sequence before being fed to the network.
Retrieval-augmented Generation with bridge model. Given the adapted passages from the bridge model as context, $(d^{\text{bdr.}}_{j})_{j=1}^{n}$, we concatenate them with the input $x$, and fed the resulting long sequence into the LLM to obtain the output for downstream tasks.
Figure 3: Illustrating the training of BGM: Step 1: First, we prepare the silver passage sequence for supervised learning (SL) through a greedy search on the retrieved passages. Step 2: This silver passage sequence is then used for the supervised training of the bridge model. The bridge model is a seq2seq model that takes the query, passage ID, and passage content as input and outputs the passage IDs. Step 3: Finally, the trained bridge model is treated as a policy model and is further trained using reinforcement learning
@@ 4 Training the Bridge Model
In Eq. 3, we format the bridge model as seq2seq model. However, it is challenging to effectively train it as there is no ground truth retrieved passages and LLM is assumed to be API without gradient. To address the issue, we propose to chain supervised learning (SL) and reinforcement learning (RL), where SL aims to reduce the search space of RL and provides a reasonably good initial model that does ranking and selection. Meanwhile, RL aims to optimize the policy model, i.e., bridge model, against the reward.
4.1 Supervised Learning
To conduct SL, the ground-truth sequence is required for each query. In a typical RAG, the ground truth relevance passage for the query is not provided, let alone the ground-truth sequence. To address this, we propose to synthesis silver passage sequence by selecting only the useful passage. This is done by greedy search that incrementally selects the next passage that can maximize the downstream task performance.
Synthesising silver passage sequence using greedy search. Assume that the downstream task performance for a given passage sequence of any length, used as context, is denoted as $R(\cdot)$. The empty passage sequence, i.e., no retrieved passage, denoted as $d^{\text{retr.}}_{\varnothing}$, has downstream task performance $R(d^{\text{retr.}}_{\varnothing})$. Starting from $d^{\text{retr.}}_{\varnothing}$ and empty silver passage sequence $d^{\text{silv.}}$, we try to add (i.e., concatenate) the candidate passage that can improve most to the silver passage sequence, until no improvement can be made. Algorithm 1 shows the pseudo-code for synthesising silver passage sequence.
Input: $(d^{\text{bdr.}}_{j})_{j=1}^{n}$, $R(\cdot)$
@@ 4.1 Supervised Learning
1 $d^{\text{silv.}}=()$;
2 $silv=R(d^{\text{retr.}}_{\varnothing})$;
3 $prev=0$;
4 $best=\varnothing$;

### 20 Feb 2024 v2

@@ Abstract
Large Language Models (LLMs) have demonstrated superior results across a wide range of tasks, and Retrieval-augmented Generation (RAG) is an effective way to enhance the performance by locating relevant information and placing it into the context window of the LLM. However, the relationship between retrievers and LLMs in a RAG is still under-investigated. Most existing work treats the retriever and the LLM as independent components and leaves a gap between retrieving human-“friendly” information and assembling a LLM-“friendly” context. In this work, we examine a novel bridge mechanism. We validate the ranking and selection assumptions of retrievers in the context of RAG and propose a framework that chains together supervised and reinforcement learning to train a bridge model that optimizes the connection between the retriever and the LLM. Empirical results demonstrate the effectiveness of our method in both question-answering and personalized generation tasks.
@@ 1 Introduction
Large language models (LLMs) such as GPT-4 openai2023gpt4 and PaLM 2 Anil et al. (2023), have demonstrated impressive performance on a wide variety of tasks. Retrieval-augmented generation (RAG), which retrieves knowledge items from an external data source and puts it into the context window of LLMs, has produced significantly enhanced results in many NLP tasks Khandelwal et al. (2020); Borgeaud et al. (2022); Izacard et al. (2022); Yasunaga et al. (2023).
@@
Figure 1: We observe a preference gap when alternating the ranking and selection of information in RAG. Experiments are conducted with retrieving passages using GTR Ni et al. (2021) and using top K of them as additional context for a frozen Palm2-S LLM. Different colors indicate different datasets (detailed in Sec. 5.1) and the Y-axis shows the relative percentage. Alternating the selection (Top-1) of information significantly affects (either positively or negatively) the LLM’s performance, while randomizing the ranking of multiple selected items (Top-5) does not have a comparable impact (the metrics are detailed in Sec. 5.2). Note the impact on NQ is even too small to be visible.
@@
However, most works on RAG study retrievers and LLMs separately. On one hand, most retrievers are designed to be human-friendly, usually based on the general belief in classic information retrieval literature that ranking is paramount, as humans typically read from top to bottom robertson1977probability. On the other hand, LLMs exhibit preferences different from humans and yield accurate results only when the information in the prompt aligns with these preferences. This discrepancy leads to sub-optimal design in current RAG systems, a phenomenon we term preference gap. This gap manifests in various aspects. For example, the general belief in ranking may not align with LLM’s preferences due to the self-attention mechanism of Transformers, which can focus on any token regardless of its position. Another aspect is selection; while humans can easily disregard irrelevant information in a context, it has been shown that LLMs are highly sensitive to irrelevant content Shi et al. (2023a). There likely exist more aspects that further diverge the LLM’s preference from that of humans, e.g., repetition. Repeated information is generally considered detrimental for retrieval systems Xia et al. (2017), but it may be useful to LLMs for weighting the relevance of context items.
@@
We empirically investigate this preference gap, focusing specifically on ranking and selection. As shown in Fig. 1, when we randomize the ordering of top-5 retrieved items (in our case, passages), the performance of RAG only varies by around 1%  [^1] We note that this finding is different from Liu et al. (2023), where a “loss in the middle” phenomenon is observed, meaning that the LLM is better at using the relevant information at the beginning or end of its input context. This is probably due to different numbers of retrieved items. Liu et al. (2023) used 20 documents, whereas we use 5 passages. Regardless, both findings indicate that there exists a preference gap between the retrievers and LLMs, which we aim to bridge.. However, the variation exceeds 5% when the LLM is only presented with the top-1 passage under each order (therefore it encounters different selections of information). This indicates the general belief in ranking does not always apply to LLMs, and that the selection of information could be more crucial. This finding confirms the existence of the preference gap between retrievers and LLMs, and it is critical to bridge this preference gap to enhance the performance of RAG. To the best of our knowledge, this is a novel insight that may guide future designs of RAG systems.
Existing work has tried to finetune the LLMs to align with the retriever or adjust the retriever to align with the LLM. However, finetuning LLMs, especially at the scale of GPT-4 or Palm 2, is often expensive. Similarly, it is difficult to adjust production-level retrievers such as Google or Bing. Even when the retriever is adjustable, existing efforts often focus on re-ranking the retrieved results and fail to address other aspects of preference such as selection or repetition. Instead, we propose a novel and practical framework called BGM (Bridging the Gap between retrievers and LLMs), which keeps the retriever and LLM fixed and trains a bridge model in between. The bridge model aims to transform the retrieved information into a format that LLMs prefers and can effectively work with.
Without loss of generality, we structure the bridge model as a sequence-to-sequence (seq2seq) model, which allows dynamically selecting items, re-ranking them, and potentially broader operations like repeating some of them in the retrieval-augmented prompt. Training such a bridge model is challenging as there are usually no ground truth labels on ideal item sequences for retrieval augmentation. Existing work has tried to derive supervisory signals for ranking from RAG’s downstream task performance, such as perplexity distillation  Izacard et al. (2022). Nevertheless, these methods only provide pointwise supervisory signals for each item independently. Directly applying the same idea to obtain sequential supervision is infeasible, since it would require feeding all possible item sequences into the LLM to obtain perplexity. We developed a greedy search approach to solve this problem (Sec. 4.1). Moreover, we find sequential supervision can be too sparse to effectively train such a seq2seq model (Table 5). To address this issue, we employed reinforcement learning (RL) on the SL trained bridge model, regarding the downstream task performance as the reward and the bridge model as a policy model. Chaining SL and RL provides increased supervision from the downstream task. It also offers the flexibility to explore more advanced strategies, such as repetition, in forming the optimal passage sequence.
@@
Our experiments reveal that BGM can enhance the performance of various downstream tasks, such as Question Answering (QA) and personalized generation, across a spectrum of datasets, from public QA and amazon reviews to private email conversations. Notably, the modified passages retrieved by BGM surpass the performance of strong retrievers and baseline reranking models. This underscores the significance and promise of the “bridge” approach in the realm of RAG. In summary, our contributions can be summarized as follows:
@@
  • We empirically establish the existence of the preference gap between retrievers and LLMs, and introduce BGM to address this gap.
@@
  • We evaluate BGM with diverse tasks, including QA and text generation, with publicly available and personalized datasets. The evaluation underscores the effectiveness of BGM in bridging the preference gap and improving RAG performance in downstream tasks.
Figure 2: Differing from previous RAGs that update LLMs, retrievers, or both, BGM connects a frozen LLM and a frozen retriever through a lightweight bridge model which adapts the retrieved information to the LLM’s preference. This makes BGM applicable to “large” LMs and off-the-shelf retrievers.
@@ 2 Related Work
Retrieval-augmented Generation (RAG). Augmenting LLMs with relevant information retrieved from various knowledge sources is proven effective across numerous NLP tasks, including language modeling Borgeaud et al. (2022); Khandelwal et al. (2020); Shi et al. (2023b), question answering Lewis et al. (2020); Izacard et al. (2022); de Jong et al. (2023); De Jong et al. (2023); Shi et al. (2023b); Guu et al. (2020); Izacard and Grave (2020); Xu et al. (2023), fact versification Lewis et al. (2020) and text generation Lewis et al. (2020). Specifically, RAG utilizes input as a query and comprises two main components: (1) a retriever retrieves a set of items from a side corpus. Particular items may vary across different tasks, including documents, passages, or even tokens. In this study, we focus on retrieving passages; and (2) a LLM incorporates the retrieved items, as additional information in the input context, and makes final predictions.
A fundamental question in this process arises regarding the disparate preferences between LLMs and retrievers, as LLMs performing optimally only when their preferences are satisfied. Bridging the preference gap is crucial. Depending on which components are subject to updates, this challenge can be categorized into three families.
Finetuning retrievers and LLMs jointly. This is the most widely used setting of RAG Izacard et al. (2022); Khandelwal et al. (2020); Wu et al. (2022); Guu et al. (2020); Lewis et al. (2020). However, most prior work is based on relative small LMs ($<$ 1B). For example, Altas Izacard et al. (2022) finetunes LLM (T5 Raffel et al. (2020a)) and retriever (Contriever Izacard et al. (2021)) jointly by leveraging the LLM to provide supervisory signal to train the retriever. RAG Lewis et al. (2020) uses a tunnable query encoder and DPR Karpukhin et al. (2020) as retriever, BART as LLM, and design an end-to-end schema to train the query encoder and the LLM.
@@
Finetuning retrievers only. Although above systems have shown improves results, they are not always applicable in practice. Many LLMs (especially larger ones) are only available as black-box APIs. Given this constraint, a natural approach is to only update the retrievers to ensure they can retrieve passages that are more compatible with LLMs. REPLUG Shi et al. (2023b) adapts a similar idea as Atlas but fix the LM. RECOMP Xu et al. (2023) trains a compressors to summarize the retrieved document from retriever. However, this family of models is incapable of performing any sample-level selection and can only choose top passages by setting a fixed threshold.
Unlike the existing approaches, BGM does not fine-tune the LLM or the retriever and instead employs a bridge model in between (Fig. 2).
RL for information retrieval. Before the LLM era, RL has been used in information retrieval (IR) Xia et al. (2017); Wei et al. (2017); Zeng et al. (2018); Xu et al. (2020). The core approach was to frame the IR problem as a Markov Decision Process and apply an RL algorithm to solve it. Typically, an IR task would be structured to determine which document to select for a ranking position, using ranking metrics such as DCG as the reward. None of these existing studies explore the application of RL in the context of RAG.
In the LLM era, RL has been used in query rewriting for retrieval Wu et al. (2021); Nogueira and Cho (2017); Adolphs et al. (2021), where a black-box retriever is assumed. This is a different problem from RAG. Bacciu et al. (2023) suggest using RL to fine-tune retriever in the RAG context in their opinion paper, not supported by experiments. Their work does not recognize the importance of bridging the gap between retrievers and LLMs, nor does it specify what the bridge model should be.
@@ 3 Problem Formulation
Retriever. Given an input $x$, the retriever aims to retrieve a ranked list of passages from a corpus $D=\{d_{i}\}_{i=1}^{m}$ that are relevant to $x$. In this work, we assume a typical scenario of employing a frozen dense retriever. Typically, a dual encoder architecture is applied, where an encoder is used to encode both the input context $x$ and the passage $d$. Specifically, the encoder maps each passage to an embedding $\bm{E}(d)$. The similarity between input and passage embedding is computed by their cosine similarity,
@@
The top-k passages that have the highest similarity scores when compared with the input $x$ are retrieved in this step,
@@
Bridge Model for RAG. The retrieved top-K passages provide richer information about the original input/query $x$ to help the LLM to make a better prediction on downstream tasks. A bridge model $\bm{B}$ adapts the retrieved passages to a sequence of passages that is LLM-friendly. As mentioned in the Sec. 1, the bridge model is a seq2seq model. It takes all the retrieved passages $(d^{\text{retr.}}_{j})_{j=1}^{k}$ as well as the query $x$ as input, and outputs the adapted passages $(d^{\text{bdr.}}_{j})_{j=1}^{n}$,
$(d^{\text{bdr.}}_{j})_{j=1}^{n}=\bm{B}\left(x,(d^{\text{retr.}}_{j})_{j=1}^{k}\right).$ (3)
This formulation is general enough as the seq2seq model automatically considers ranking by generating the next token based on the preceding one, selection by placing the end-of-sentence token in the appropriate position, and repetion by generating the same passage ID (as explained in the following paragraph). Note that $n$ may be smaller or larger than $k$ due to selection and repetition.
Before concatenating the query and passages as bridge model’s input, we prepend each passage with a unique sentinel token as its passage ID, e.g., $[\text{query}][\text{id}_{1}]d^{\text{retr.}}_{1}[\text{id}_{2}]d^{\text{retr.}}_{2}$. In this way, the model only needs to generate the passage IDs instead of the actual passage content, which is much more efficient and avoids making unfaithful changes to the retrieved passages. We then convert the obtained passage IDs to the corresponding passages for downstream processing.
Retrieval-augmented generation with bridge. We concatenate adapted passages from the bridge model, $(d^{\text{bdr.}}_{j})_{j=1}^{n}$, with the input $x$, and fed the resulting long sequence into the LLM as context to obtain the output for downstream tasks.
Figure 3: Illustrating the training of BGM: Step 1: First, we prepare the silver passage sequence for supervised learning (SL) through a greedy search on the retrieved passages. Step 2: These silver passage sequences are then used for the supervised training of the bridge model. The bridge model is a seq2seq model that takes the query and passages with prepended passage IDs as input and outputs the passage IDs. Step 3: Finally, the SL-trained bridge model is treated as a policy model and is further trained using reinforcement learning.
@@ 4 Training the Bridge Model
In Eq. 3, we format the bridge model as seq2seq model. As mentioned in Sec. 1, its training is challenging due to the lack of supervision for passage sequences, the infeasibility of directly applying existing methods that provide only point-wise supervision, and the sparsity of sequential supervision. For effective training, we propose to chain supervised learning (SL) and reinforcement learning (RL), where SL aims to reduce the search space of RL and provides a reasonably good initial model that does ranking and selection. RL aims to optimize the policy model, i.e., bridge model, for the downstream task. Fig. 3 shows an overview.
Input: $(d^{\text{retr.}}_{j})_{j=1}^{n}$, $R(\cdot)$
@@
1 $d^{\text{silv.}}\leftarrow()$;
2 $R^{\text{silv.}}\leftarrow R(d^{\text{silv.}})$;
3 $R^{\text{best}}\leftarrow-\infty$;